# Interconnect-Aware Logic Resynthesis for Multi-Die FPGAs

### v1

@@ 4.3. Impact on Inter-Die Connectivity
We evaluate the effectiveness of ResynMD on pure combinational logic benchmarks from the EPFL benchmark suite. The input netlists are first partitioned using KaHyPar (Schlag et al., 2022) with equal node weights, without considering timing criticality. Subsequently, Algorithm 1 is applied to perform interconnect-aware logic resynthesis. As illustrated in Figure 9, the inter-die connectivity before and after interconnect-aware resynthesis highlights the effectiveness of the proposed optimization, where plenty of inter-die connections (shown in red) are eliminated.
@@ 4.5.1. Inter-Die Connectivity Reduction
Table 6 reports the impact of ResynMD on inter-die connectivity under the 2-die configuration. Across the MCNC (Yang, 1991) benchmarks, the proposed method reduces inter-die fanout edges by 12.29% on average, while the net-level reduction is 1.65%. The discrepancy between edge-level and net-level improvements is expected. Multiple cross-die sinks may share a single SLL track, and removing individual fanout edges does not necessarily eliminate the occupied SLL channel. Nevertheless, reducing inter-die fanout edges lowers SLL loading and simplifies inter-die connectivity at the structural level.

### v2

@@ 4.3. Impact on Inter-Die Connectivity
We evaluate the effectiveness of ResynMD on pure combinational logic benchmarks from the EPFL benchmark suite. The input netlists are first partitioned using KaHyPar (KaHyPar, ) with equal node weights, without considering timing criticality. Subsequently, Algorithm 1 is applied to perform interconnect-aware logic resynthesis. As illustrated in Figure 9, the inter-die connectivity before and after interconnect-aware resynthesis highlights the effectiveness of the proposed optimization, where plenty of inter-die connections (shown in red) are eliminated.
@@ 4.5.1. Inter-Die Connectivity Reduction
Table 6 reports the impact of ResynMD on inter-die connectivity under the 2-die configuration. Across the MCNC (yang1991logic, ) benchmarks, the proposed method reduces inter-die fanout edges by 12.29% on average, while the net-level reduction is 1.65%. The discrepancy between edge-level and net-level improvements is expected. Multiple cross-die sinks may share a single SLL track, and removing individual fanout edges does not necessarily eliminate the occupied SLL channel. Nevertheless, reducing inter-die fanout edges lowers SLL loading and simplifies inter-die connectivity at the structural level.